OpenStudy (anonymous):

can you help me solve x/2+x/3=7

OpenStudy (anonymous):

are those division or fractions symbols??

OpenStudy (anonymous):

fraction symbols

OpenStudy (anonymous):

kk

zepdrix (zepdrix):

\[\Large \color{royalblue}{\text{Welcome To OpenStudy! :)}}\] Is this how your problem appears?\[\Large \frac{x}{2}+\frac{x}{3}=7\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

First we remove the denominator by multiplying all terms by the LCD 6 That will give you" 3x + 2x = 42 Continue from there, answer will be a fraction

OpenStudy (anonymous):

Thank you for all the help I received

OpenStudy (anonymous):

np